OpenStudy (anonymous):

lim (x+1)sin(4/x) x-> infinity

OpenStudy (alfie):

\[\lim_{x->\infty} (x+1)\sin(\frac{4}{x})\] You divide and multiply for sin(4/x) \[\lim_{x->\infty} \frac{4x+4}{x} \frac{\sin(\frac{4}{x})}{\frac{4}{x}}\] The second goes to 0 and usuing senx/x you know it goes to one. You factorize four up there, getting... \[1 * \lim_{x->\infty} \frac{4(x+1)}{x} \] 4 is a constant, you take it outside the limit. Everything goes at one, you have the four outside, so you have \[\lim_{x->\infty} (x+1)\sin(\frac{4}{x}) = 4\]

OpenStudy (anonymous):

Thanks!

OpenStudy (alfie):

You're welcome :)